A juggler throws a ball from height of 0.950 m with a vertical velocity of +3.50 m/s and misses it on the way down. What is its velocity when it hits the ground?
1
Expert's answer
2010-12-23T03:12:52-0500
According to the equations of the motion:
h = h0 +v0t - gt2/2;
v = v0 - gt.


h0 = 0.95
v0 = 3.5 m/s
h = 0

0 = 0.95 +3.5t - 10t2/2
5t2 - 3.5t - 0.95 = 0
D = 3.52 + 5*0.95*4 = 31.25
t1 = (3.5+5.6) / 10 = 0.91
t2 = -0.21 - out of sense.

v = 3.5 - 0.91*10 = -5.6 m/s
Answer:
The velocity would be equal to 5.6 m/s and its vector is directed downward.